OpenStudy (anonymous):

calc 3 help plz use lagrange multipliers to find the max and min values of f(x,y,z)=2x+y-2z subject to the constraint x^2+y^2+z^2=4 and find the points at which these extreme values occur.

OpenStudy (perl):

this isnt too bad, check on pauls online notes

OpenStudy (perl):

would you like to go through it?

OpenStudy (anonymous):

uhm sure.

OpenStudy (anonymous):

Should I just say what I did?

OpenStudy (perl):

yeah

OpenStudy (anonymous):

found gradient of F and lamda gradient of G

OpenStudy (anonymous):

put like terms together form the 2 functions, the i components together, j components and z components

OpenStudy (perl):

ok lets set it up

OpenStudy (anonymous):

uhm I did gradient F = lamda gradient G

OpenStudy (perl):

solve the system <fx, fy,fz> = lambda <gx,gy,gz> g(x,y,z) = k

OpenStudy (anonymous):

found x=1/lamda, y=1/2lamda and z = -1/lamda

OpenStudy (perl):

one sec

OpenStudy (perl):

<2,1,-2> = L <2x, 2y, 2z> 2 = 2Lx, 1 = 2Ly , -2 = 2Lz

OpenStudy (anonymous):

Yes I got that, then I solved for x, y and z

OpenStudy (perl):

ok so we need to solve the system of equations 2 = 2Lx, 1 = 2Ly , -2 = 2Lz , x^2+y^2+z^2=4

OpenStudy (perl):

so I get 1/x = 1/(2y)= -1/z ,

OpenStudy (anonymous):

Yes, I plugged in the values I got, and solved for lamda

OpenStudy (anonymous):

yes

OpenStudy (perl):

so x/2 = y, and z = -x , we can plug that into g(x) x^2 + (x/2)^2 + (-x)^2 = 4

OpenStudy (anonymous):

Would it be right if I just plugged in the x, y, and z values?

OpenStudy (perl):

x^2 +(x/2)^2 + (-x)^2 = 4 , solve that

OpenStudy (anonymous):

Like rather than subbing it in terms of X, would I be able to just put in 1/L for x and 1/2L for y etc?

OpenStudy (perl):

you want to solve for x, not for L

OpenStudy (anonymous):

Ok, but before I solve it

OpenStudy (perl):

L is just an auxiliary thing

OpenStudy (anonymous):

why isn't it x^2+(2x)^2+(-x)^2=4?

OpenStudy (perl):

do you agree that 2 = 2Lx, 1 = 2Ly , -2 = 2Lz , x^2+y^2+z^2=4

OpenStudy (anonymous):

Yes

OpenStudy (perl):

L = 1/x , L = 1/(2y ) , L = -1/z now set them equal to each other, since they all equal to L

OpenStudy (anonymous):

oohhh, that changes everything, I solved for the variables, not L

OpenStudy (perl):

basically we want to get rid of the L. it is just a helper, we dont care about L actually

OpenStudy (perl):

it helps to turn this into a one variable problem,

OpenStudy (anonymous):

So after I solve for L

OpenStudy (perl):

actually you can solve for L if you want

OpenStudy (anonymous):

I just erased my work =[[

OpenStudy (perl):

no problem

OpenStudy (perl):

in this problem, either way works actually

OpenStudy (anonymous):

after you solved L = 1/x, 1/2y, -1/z what do you do?

OpenStudy (anonymous):

so I got x = positive or negative 4/3

OpenStudy (anonymous):

hhhmm

OpenStudy (perl):

do you agree x^2 +(x/2)^2 + (-x)^2 = 4

OpenStudy (anonymous):

yeah that what i got

OpenStudy (anonymous):

x^2+(x^2/4)+x^2=4

OpenStudy (perl):

x^2 + x^2/4 + x^2 = 4 9/4 *x^2 = 4

OpenStudy (anonymous):

its 16/9

OpenStudy (anonymous):

x^2= 16/9

OpenStudy (perl):

right, woops

OpenStudy (perl):

ok , i agree x = + - 4/3

OpenStudy (anonymous):

Yes, I almost erased it again aahhaha

OpenStudy (anonymous):

from there you plug x back into L = 1/L correct?

OpenStudy (anonymous):

then solve for L, and plug L back into the other ones?

OpenStudy (perl):

hmmm, we plug this x into f(x,y,z) first find y and z though

OpenStudy (perl):

we have two solutions for x , (4/3 , , ) (-4/3 , , )

OpenStudy (anonymous):

for my Y i got positive and negative 2/3

OpenStudy (anonymous):

and x = -z so positive and negative 4/3?

OpenStudy (perl):

yes

OpenStudy (anonymous):

Alright, then plug in EVERY combination of those points, back into the F function, find which one gives you the highest and lowest value, and that's the max/min and the points are the coordinates for which they occur?

OpenStudy (perl):

so we test (4/3, 2/3, -4/3) and (-4/3, -2/3, 4/3)

OpenStudy (perl):

right

OpenStudy (anonymous):

I got 8 different points

OpenStudy (anonymous):

but the points you just said.... are the max/min points ahah, how'd you figure them out so quick?

OpenStudy (perl):

8 different points?

OpenStudy (perl):

since x = 4/3, -4/3 that means there will be two points, remember y and z depend on x here because of our work above

OpenStudy (perl):

two points (x,y,z) to plug into f(x,y,z) , one will be greater than the other, so that tells you which is min which is max

OpenStudy (anonymous):

oh you're right, x = -z

OpenStudy (perl):

you could solve this by using L, you were right earlier